OpenStudy (anonymous):

what is the value of 0! 9! / (9-2)!

jimthompson5910 (jim_thompson5910):

The exclamation mark is the factorial symbol It means that you multiply that number and every whole number smaller than it until you reach 1 So 9! = 9*8*7*6*5*4*3*2*1 8! = 8*7*6*5*4*3*2*1 7! = 7*6*5*4*3*2*1 etc etc

jimthompson5910 (jim_thompson5910):

Tell me what you get

OpenStudy (ikram002p):

0!=1 9!=9*8*7! (9-2)!=7!

OpenStudy (anonymous):

1(362880) / 5040 ?

jimthompson5910 (jim_thompson5910):

good, now reduce

OpenStudy (anonymous):

72

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

so is that the answer?

jimthompson5910 (jim_thompson5910):

yes it is

OpenStudy (anonymous):

Thank you!

OpenStudy (ikram002p):

ugh why dnt u just do this \(\Huge \frac{0! 9!}{(9-2)!}\) \(\Huge \frac{1×9×8× 7!}{(7)!}= 9×8 \)